# Efficient Fourier Filtering Network with Contrastive Learning for UAV-based Unaligned Bi-modal Salient Object Detection

### v2

@@ IV-E Ablation Study
To assess the contribution of each key component, namely the backbone network, SCAL, and SAF, in our proposed AlignSal, we conducted ablation experiments on the UAV RGB-T 2400 dataset.
1) Effectiveness of backbone network: Owing to its efficient and powerful feature-generating capability, we chose the FFT-based CDFFormer-S18 [57] as the backbone for our AlignSal. To verify its effectiveness, we selected three other FFT-based backbones to replace CDFFormer-S18, including GFNet-H-S [58], AFNO [59], and DFFormer-S18 [57]. While AFNO had the same network depth as CDFFormer-S18, the comparison in Table IV showed that CDFFormer-S18 has significant advantages over other models.
To further assess the compatibility of AlignSal with spatial domain-based backbone networks, we conducted experiments using ResNet-50 [75] and Swin Transformer Tiny (Swin-T) [76] as backbone networks. Table IV demonstrates that AlignSal can be effectively integrated with these spatial domain-based backbone networks. However, our findings also show that their performance is weaker than FFT-based backbone networks when computational resources are limited.
@@
Figure 6: Visual comparison of the model with and without (w/o) the key components on the UAV RGB-T 2400 dataset. GT represents ground truth.
2) Effectiveness of SCAL: To assess the specific contribution of SCAL, we conducted ablation experiments from various perspectives. First, we removed SCAL and presented the results in Table IV, Fig. 5, and Fig. 6. The $Em$, $Sm$, $wF_{\beta}$, $F_{\beta}$, and $\mathcal{M}$ of the model without SCAL, as shown in Table IV, decreased significantly by 0.7%, 1.1%, 4.6%, 3.4%, and 14.5%, respectively. Missing the semantic-level alignment resulted in inconsistencies in the salient regions highlighted by the two modalities, adversely affecting the final predictions. Specifically, comparing (b) and (c) in Fig. 5, the salient regions in the RGB feature maps of the model without SCAL were less distinct and exhibited noticeable drifts. In contrast, the model with SCAL not only resolved these issues, but also, as illustrated by comparing (d) and (e) in Fig. 5, enabled the thermal features to benefit from the RGB features, better distinguishing salient regions from the background. Additionally, SCAL encouraged the model to reuse low-level features more efficiently, improving feature representation and reducing the burden of subsequent pixel-level alignment. In the results of the model without SCAL shown in the $1^{st}$ row of Fig. 6, the leftmost ground area (i.e., the area in red box) was incorrectly predicted as salient due to its highly distinctive features in the thermal modality. Similarly, the rightmost bike in the $2^{nd}$ row was misclassified as the background.
Second, to further validate the performance of SCAL in enhancing the model to handle unaligned data, we incorporated SCAL into the training phase of representative aligned BSOD models (i.e., MIDD [13], MoADNet [15], LSNet [46], MGAI [12], and CATNet [19]) on the UAV RGB-T 2400 dataset. In this experiment, we set the values of $k$ and $t$ of SCAL to 3 and 0.5, respectively, without fine-tuning. As seen in Table VII, the performance gains from SCAL were substantial. The $Em$, $Sm$, $wF_{\beta}$, $F_{\beta}$, and $\mathcal{M}$ improved by an average of 2.0%, 2.1%, 7.0%, 4.7%, and 16.2%, respectively.
@@
TABLE V: Quantitative Comparison Between Our AlignSal and Eleven State-of-the-Art BSOD Models on the Unaligned-VT821, Unaligned-VT1000, and Unaligned-VT5000 Datasets [53]. The Best Results are Labeled Red and the Second Best Results are Labeled Blue. We Tested the FPS of Different Models on an NVIDIA RTX4060 Laptop GPU.
@@
CATNet23 [19] 0.8934 0.8596 0.7719 0.7899 0.0404 0.9289 0.9203 0.8855 0.8787 0.0224 0.9137 0.8705 0.8065 0.8176 0.0352 262.6M 341.8G 19.7
PICRNet23 [52] 0.8339 0.7917 0.6644 0.6918 0.0696 0.8545 0.8285 0.7374 0.7536 0.0670 0.8084 0.7610 0.6243 0.6517 0.0920 112.0M 27.1G 30.3
LAFB24 [23] 0.8917 0.8588 0.7772 0.7959 0.0502 0.9420 0.9264 0.8961 0.8985 0.0206 0.9225 0.8836 0.8233 0.8400 0.0337 453.0M 139.7G 29.3
DCNet22 [53] 0.9084 0.8599 0.7986 0.8262 0.0364 0.9432 0.9149 0.8891 0.9015 0.0233 0.9075 0.8536 0.7899 0.8237 0.0409 24.1M 207.2G 14.5
@@
TABLE VI: Quantitative Comparison Between Our AlignSal and Lightweight (L) and Heavyweight (H) BSOD Models on the VT821 [36], VT1000 [66], and VT5000 [11] Datasets. The Best Results are Labeled Red and the Second Best Results are Labeled Blue. We Tested the FPS of Different Models on an NVIDIA RTX4060 Laptop GPU.
@@
TABLE VII: Quantitative Comparison of Representatives BSOD Models Before and After Being Equipped with Semantic Contrastive Alignment Loss (SCAL) on the UAV RGB-T 2400 Dataset [1]. * Stands for the BSOD Model with SCAL. Subscripted Numbers Indicate Performance Improvement or Degradation Compared to the Original Model.
@@
MIDD 0.8298-${}_{\emph{{\color[rgb]{0,0,1}-0.0035}}}$ 0.8593${}_{\emph{{\color[rgb]{1,0,0}+0.0069}}}$ 0.7336${}_{\emph{{\color[rgb]{1,0,0}+0.0168}}}$ 0.6581${}_{\emph{{\color[rgb]{1,0,0}+0.0198}}}$ 0.0103-${}_{\emph{{\color[rgb]{1,0,0}-0.0011}}}$
MoADNet 0.4744${}_{\emph{{\color[rgb]{1,0,0}+0.0117}}}$ 0.5019${}_{\emph{{\color[rgb]{1,0,0}+0.0036}}}$ 0.0724${}_{\emph{{\color[rgb]{1,0,0}+0.0133}}}$ 0.1147${}_{\emph{{\color[rgb]{1,0,0}+0.0162}}}$ 0.0332-${}_{\emph{{\color[rgb]{1,0,0}-0.0038}}}$
LSNet 0.8948${}_{\emph{{\color[rgb]{1,0,0}+0.0021}}}$ 0.7189${}_{\emph{{\color[rgb]{1,0,0}+0.0242}}}$ 0.3578${}_{\emph{{\color[rgb]{1,0,0}+0.0405}}}$ 0.6795-${}_{\emph{{\color[rgb]{0,0,1}-0.0001}}}$ 0.0330-${}_{\emph{{\color[rgb]{1,0,0}-0.0081}}}$
MGAI 0.8629${}_{\emph{{\color[rgb]{1,0,0}+0.0237}}}$ 0.8471${}_{\emph{{\color[rgb]{1,0,0}+0.0175}}}$ 0.7212${}_{\emph{{\color[rgb]{1,0,0}+0.0415}}}$ 0.6444${}_{\emph{{\color[rgb]{1,0,0}+0.0386}}}$ 0.0113-${}_{\emph{{\color[rgb]{1,0,0}-0.0010}}}$
CATNet 0.7868${}_{\emph{{\color[rgb]{1,0,0}+0.0448}}}$ 0.8412${}_{\emph{{\color[rgb]{1,0,0}+0.0275}}}$ 0.7024${}_{\emph{{\color[rgb]{1,0,0}+0.0693}}}$ 0.5698${}_{\emph{{\color[rgb]{1,0,0}+0.0509}}}$ 0.0155-${}_{\emph{{\color[rgb]{1,0,0}-0.0027}}}$
significantly enhancing their performance in handling unaligned bi-modal images without taking up additional computational resources during inference.
3) Effectiveness of SAF: To evaluate the effectiveness of SAF, we replaced it with element-wise addition for bi-modal fusion, referred to as “w/o SAF”. As shown in the $5^{th}$ row of Table IV, the performance of AlignSal without SAF significantly declined. This finding was further supported by Fig. 5 and 6. In Fig. 5 (g), the model without SAF failed to align RGB and thermal features in the fused features and did not effectively integrate the bi-modal complementary information, resulting in confused and overlapping salient regions. These issues were particularly noticeable in high-resolution fused feature maps. In Fig. 6, the model without SAF failed to capture detailed features (e.g., the small human figures in the $1^{st}$ row) and to discriminate inter-modal interference information (e.g., the statue in the $3^{rd}$ row). Moreover, we observed that the full model maintained consistent Params with only 0.1G higher FLOPs compared to the model without SAF. This indicates that the complexity of our proposed SAF is comparable to that of element-wise addition.
@@ IV-F Generalization Experiment
To further validate the model’s generalization, we compared our AlignSal with existing state-of-the-art bi-modal dense prediction models across three different tasks: weakly aligned bi-modal salient object detection (BSOD), aligned BSOD, and remote sensing change detection (RSCD). All prediction maps used for
TABLE VIII: Quantitative Comparison Between Our AlignSal and Three Existing RSCD Models on the LEVIR-CD+ Dataset [67]. The Best Results are Labeled Red.
@@
FC-EF18 [60] 71.77 69.12 97.54 70.42 54.34 1.4M 14.1G
FC-Siam-Diff18 [60] 74.02 81.49 98.26 77.57 63.36 1.4M 18.7G
SNUNet22 [77] 78.73 71.07 97.83 74.70 59.62 10.2M 176.4G
HANet23 [78] 75.53 79.70 98.22 77.56 63.34 2.6M 70.7G
@@
1) Weakly Aligned BSOD: We first compared our AlignSal with existing BSOD models, including MIDD [13], MoADNet [15], DCNet [53], OSRNet [17], LSNet [46], TNet [18], MGAI [12], CATNet [19], PICRNet [52], LAFB [23], and MROS [1] on three unaligned datasets, namely unaligned-VT821, unaligned-VT1000, and unaligned-VT5000 [53].
As shown in Table V, AlignSal achieved a significant lead with lower computational cost and faster inference speed compared to the current top-performing models, such as MROS [1], DCNet [53], and LAFB [23]. These comprehensive experiments confirmed that AlignSal, with our proposed key components, can generalize well to different unaligned bi-modal datasets.
2) Aligned BSOD: We then compared our AlignSal with three lightweight BSOD models, namely CSRNet [14], OSRNet [17], and LSNet [46], and eight heavyweight BSOD models, namely MIDD [13], DCNet [53], SwinNet [16], TNet [18], HRTNet [20], MGAI [12], WaveNet [22], and LAFB [23] on the VT821 [36], VT1000 [66], and VT5000 [11] datasets.
@@
3) RSCD: To evaluate the generalizability of our AlignSal in high-altitude scenarios, we compared it with three existing RSCD models, including FC [60], SNUNet [77], and HANet [78] on the LEVIR-CD+ dataset [67].
@@ V Conclusions
In this paper, we examined the relationship between unaligned modalities and proposed AlignSal, which achieves both real-time performance and high accuracy for UAV-based unaligned BSOD. To this end, we first designed SCAL to align bi-modal semantic features in a contrastive manner and enhance inter-modal information exchange. Notably, SCAL not only works on our proposed AlignSal, but also improves the performance of various existing aligned BSOD models on the UAV RGB-T 2400 dataset. Second, we proposed the FFT-based SAF to align slightly offset features in multiple dimensions and facilitate bi-modal fusion by efficiently acquiring global relevance. Extensive experiments demonstrated that AlignSal achieves significantly faster inference speed and superior performance compared to existing models on the UAV RGB-T 2400 dataset, three weakly aligned BSOD datasets, three aligned BSOD datasets, and one remote sensing change detection dataset. These advantages make AlignSal suitable for deployment on UAVs in practice, enabling its use in industrial and real-life applications that require both timeliness and accuracy.

### v3

@@ IV-E Ablation Study
To assess the contribution of each key component, namely the backbone network, SCAL, and SAF, in our proposed AlignSal, we conducted ablation experiments on the AAV RGB-T 2400 dataset.
1) Effectiveness of backbone network: Owing to its efficient and powerful feature-generating capability, we chose the FFT-based CDFFormer-S18 [tatsunami2024fft] as the backbone for our AlignSal. To verify its effectiveness, we selected three other FFT-based backbones to replace CDFFormer-S18, including GFNet-H-S [rao2021global], AFNO [guibas2021adaptive], and DFFormer-S18 [tatsunami2024fft]. While AFNO had the same network depth as CDFFormer-S18, the comparison in Table IV showed that CDFFormer-S18 has significant advantages over other models.
To further assess the compatibility of AlignSal with spatial domain-based backbone networks, we conducted experiments using ResNet-50 [he2016deep] and Swin Transformer Tiny (Swin-T) [liu2021swin] as backbone networks. Table IV demonstrates that AlignSal can be effectively integrated with these spatial domain-based backbone networks. However, our findings also show that their performance is weaker than FFT-based backbone networks when computational resources are limited.
@@
Figure 6: Visual comparison of the model with and without (w/o) the key components on the AAV RGB-T 2400 dataset. GT represents ground truth.
2) Effectiveness of SCAL: To assess the specific contribution of SCAL, we conducted ablation experiments from various perspectives. First, we removed SCAL and presented the results in Table IV, Fig. 5, and Fig. 6. The $Em$, SmSm, $wF_{\beta}$, $F_{\beta}$, and $\mathcal{M}$ of the model without SCAL, as shown in Table IV, decreased significantly by 0.7%, 1.1%, 4.6%, 3.4%, and 14.5%, respectively. Missing the semantic-level alignment resulted in inconsistencies in the salient regions highlighted by the two modalities, adversely affecting the final predictions. Specifically, comparing (b) and (c) in Fig. 5, the salient regions in the RGB feature maps of the model without SCAL were less distinct and exhibited noticeable drifts. In contrast, the model with SCAL not only resolved these issues, but also, as illustrated by comparing (d) and (e) in Fig. 5, enabled the thermal features to benefit from the RGB features, better distinguishing salient regions from the background. Additionally, SCAL encouraged the model to reuse low-level features more efficiently, improving feature representation and reducing the burden of subsequent pixel-level alignment. In the results of the model without SCAL shown in the $1^{st}$ row of Fig. 6, the leftmost ground area (i.e., the area in red box) was incorrectly predicted as salient due to its highly distinctive features in the thermal modality. Similarly, the rightmost bike in the $2^{nd}$ row was misclassified as the background.
Second, to further validate the performance of SCAL in enhancing the model to handle unaligned data, we incorporated SCAL into the training phase of representative aligned BSOD models (i.e., MIDD [9454273], MoADNet [jin2022moadnet], LSNet [10042233], MGAI [10003255], and CATNet [sun2023catnet]) on the AAV RGB-T 2400 dataset. In this experiment, we set the values of $k$ and $t$ of SCAL to 3 and 0.5, respectively, without fine-tuning. As seen in Table VII, the performance gains from SCAL were substantial. The $Em$, SmSm, $wF_{\beta}$, $F_{\beta}$, and $\mathcal{M}$ improved by an average of 2.0%, 2.1%, 7.0%, 4.7%, and 16.2%, respectively.
@@
TABLE V: Quantitative Comparison Between Our AlignSal and Eleven State-of-the-Art BSOD Models on the Unaligned-VT821, Unaligned-VT1000, and Unaligned-VT5000 Datasets [tu2022weakly]. The Best Results are Labeled Red and the Second Best Results are Labeled Blue. We Tested the FPS of Different Models on an NVIDIA RTX4060 Laptop GPU.
@@
CATNet23 [sun2023catnet] 0.8934 0.8596 0.7719 0.7899 0.0404 0.9289 0.9203 0.8855 0.8787 0.0224 0.9137 0.8705 0.8065 0.8176 0.0352 262.6M 341.8G 19.7
PICRNet23 [cong2023point] 0.8339 0.7917 0.6644 0.6918 0.0696 0.8545 0.8285 0.7374 0.7536 0.0670 0.8084 0.7610 0.6243 0.6517 0.0920 112.0M 27.1G 30.3
LAFB24 [wang2024learning] 0.8917 0.8588 0.7772 0.7959 0.0502 0.9420 0.9264 0.8961 0.8985 0.0206 0.9225 0.8836 0.8233 0.8400 0.0337 453.0M 139.7G 29.3
DCNet22 [tu2022weakly] 0.9084 0.8599 0.7986 0.8262 0.0364 0.9432 0.9149 0.8891 0.9015 0.0233 0.9075 0.8536 0.7899 0.8237 0.0409 24.1M 207.2G 14.5
@@
TABLE VI: Quantitative Comparison Between Our AlignSal and Lightweight (L) and Heavyweight (H) BSOD Models on the VT821 [wang2018rgb], VT1000 [tu2019rgb], and VT5000 [9767629] Datasets. The Best Results are Labeled Red and the Second Best Results are Labeled Blue. We Tested the FPS of Different Models on an NVIDIA RTX4060 Laptop GPU.
@@
TABLE VII: Quantitative Comparison of Representatives BSOD Models Before and After Being Equipped with Semantic Contrastive Alignment Loss (SCAL) on the AAV RGB-T 2400 Dataset [10315195]. * Stands for the BSOD Model with SCAL. Subscripted Numbers Indicate Performance Improvement or Degradation Compared to the Original Model.
@@
[…]
significantly enhancing their performance in handling unaligned bimodal images without taking up additional computational resources during inference.
3) Effectiveness of SAF: To evaluate the effectiveness of SAF, we replaced it with element-wise addition for bimodal fusion, referred to as “w/o SAF”. As shown in the $5^{th}$ row of Table IV, the performance of AlignSal without SAF significantly declined. This finding was further supported by Fig. 5 and 6. In Fig. 5 (g), the model without SAF failed to align RGB and thermal features in the fused features and did not effectively integrate the bimodal complementary information, resulting in confused and overlapping salient regions. These issues were particularly noticeable in high-resolution fused feature maps. In Fig. 6, the model without SAF failed to capture detailed features (e.g., the small human figures in the $1^{st}$ row) and to discriminate inter-modal interference information (e.g., the statue in the $3^{rd}$ row). Moreover, we observed that the full model maintained consistent Params with only 0.1G higher FLOPs compared to the model without SAF. This indicates that the complexity of our proposed SAF is comparable to that of element-wise addition.
@@ IV-F Generalization Experiment
To further validate the model’s generalization, we compared our AlignSal with existing state-of-the-art bimodal dense prediction models across three different tasks: weakly aligned bimodal salient object detection (BSOD), aligned BSOD, and remote sensing change detection (RSCD). All prediction maps used for
TABLE VIII: Quantitative Comparison Between Our AlignSal and Three Existing RSCD Models on the LEVIR-CD+ Dataset [chen2020spatial]. The Best Results are Labeled Red.
@@
FC-EF18 [daudt2018fully] 71.77 69.12 97.54 70.42 54.34 1.4M 14.1G
FC-Siam-Diff18 [daudt2018fully] 74.02 81.49 98.26 77.57 63.36 1.4M 18.7G
SNUNet22 [fang2021snunet] 78.73 71.07 97.83 74.70 59.62 10.2M 176.4G
HANet23 [han2023hanet] 75.53 79.70 98.22 77.56 63.34 2.6M 70.7G
@@
1) Weakly Aligned BSOD: We first compared our AlignSal with existing BSOD models, including MIDD [9454273], MoADNet [jin2022moadnet], DCNet [tu2022weakly], OSRNet [9803225], LSNet [10042233], TNet [9926193], MGAI [10003255], CATNet [sun2023catnet], PICRNet [cong2023point], LAFB [wang2024learning], and MROS [10315195] on three unaligned datasets, namely unaligned-VT821, unaligned-VT1000, and unaligned-VT5000 [tu2022weakly].
As shown in Table V, AlignSal achieved a significant lead with lower computational cost and faster inference speed compared to the current top-performing models, such as MROS [10315195], DCNet [tu2022weakly], and LAFB [wang2024learning]. These comprehensive experiments confirmed that AlignSal, with our proposed key components, can generalize well to different unaligned bimodal datasets.
2) Aligned BSOD: We then compared our AlignSal with three lightweight BSOD models, namely CSRNet [9505635], OSRNet [9803225], and LSNet [10042233], and eight heavyweight BSOD models, namely MIDD [9454273], DCNet [tu2022weakly], SwinNet [9611276], TNet [9926193], HRTNet [9869666], MGAI [10003255], WaveNet [10127616], and LAFB [wang2024learning] on the VT821 [wang2018rgb], VT1000 [tu2019rgb], and VT5000 [9767629] datasets.
@@
3) RSCD: To evaluate the generalizability of our AlignSal in high-altitude scenarios, we compared it with three existing RSCD models, including FC [daudt2018fully], SNUNet [fang2021snunet], and HANet [han2023hanet] on the LEVIR-CD+ dataset [chen2020spatial].
@@ V Conclusions
In this paper, we examined the relationship between unaligned modalities and proposed AlignSal, which achieves both real-time performance and high accuracy for AAV-based unaligned BSOD. To this end, we first designed SCAL to align bimodal semantic features in a contrastive manner and enhance inter-modal information exchange. Notably, SCAL not only works on our proposed AlignSal, but also improves the performance of various existing aligned BSOD models on the AAV RGB-T 2400 dataset. Second, we proposed the FFT-based SAF to align slightly offset features in multiple dimensions and facilitate bimodal fusion by efficiently acquiring global relevance. Extensive experiments demonstrated that AlignSal achieves significantly faster inference speed and superior performance compared to existing models on the AAV RGB-T 2400 dataset, three weakly aligned BSOD datasets, three aligned BSOD datasets, and one remote sensing change detection dataset. These advantages make AlignSal suitable for deployment on AAVs in practice, enabling its use in industrial and real-life applications that require both timeliness and accuracy.